In my textbook it says to solve the System of Linear Equations using Substitution. The problem it gives for it is:
x= 8y-4
x+8y=6
Do you know how to work this out?

Answer by mananth(16946)

x=8y-4
x+8y=6
In place of x substitute x from first equation
(8y-4)+8y=6
16y-4=6
+4 +4
16y=10
/16
y=10/16
y=5/8
plug y in first equation
x=8y-4
x=8*(5/8)-4
x=1